# Evidence Contextualization and Counterfactual Attribution for Conversational QA over Heterogeneous Data with RAG Systems

### v1

@@ Abstract.
ccs: Information systems Question answering
@@ 1. Introduction
The benchmark and all other artifacts for this work are public at https://github.com/Fraunhofer-IIS/RAGonite.
@@ 2. System overview
Backend. An overview of the RAGonite pipeline is in Fig. 1. RAGonite’s backend is split into a functional core, which handles retrieval and answer generation, and a stateful layer that persists chats into a SQLite database and provides a REST API to the frontend using FastAPI. The dependencies in the functional core include the vector database (ChromaDB), and prompt template (Jinja) and LLM libraries (gpt-4o). We use ChromaDB as our vector database for storing the contextualized evidences, and also use ChromaDB’s in-built retrieval functions. The multilingual BGE embedding model [^1] https://huggingface.co/BAAI/bge-m3 was used to embed the documents, that was found to work slightly better than text-embedding-3-small [^2] https://platform.openai.com/docs/guides/embeddings/embedding-models from OpenAI. While ChromaDB was used for its extensibility, we also explored variants like Weaviate and Milvus. We used top-k hybrid search with the BGE re-ranker [^3] https://huggingface.co/BAAI/bge-reranker-v2-m3 being used to merge $k$ dense retrieval and lexical retrieval results ($k=10$ in our experiments). We use gpt-4o as our LLM of choice inside the question completion and answer generation modules for efficiency and quality, but we also support other OpenAI, Mistral, and Llama models. All prompts can be seen inside the demo interface for transparency. A single GPU server (4x48GB NVIDIA Ada 6000 RTX, 512 GB RAM, 64 virtual cores) was used for all our experiments. All code is in Python.
@@ 2 Output: Distribution $\boldsymbol{\mathcal{A}}$ for attributing answer $a$ to each evidence $e$
Metric $\rightarrow$ Retrieval precision@1 Answer relevance
Questions $\downarrow$ / Context $\rightarrow$ NONE +TTL +HDR +BEF +AFT +ALL NONE +TTL +HDR +BEF +AFT +ALL
All questions ($600$) $0.440$ $0.493$ $0.472$ $0.448$ $0.450$ $\boldsymbol{0.523}$* $0.435$ $0.538$ $0.494$ $0.478$ $0.484$ $\boldsymbol{0.585}$*
Simple questions ($300$) $0.453$ $0.477$ $0.480$ $0.457$ $0.460$ $\boldsymbol{0.500}$* $0.492$ $0.573$ $0.560$ $0.545$ $0.538$ $\boldsymbol{0.678}$*
Complex questions ($300$) $0.427$ $0.510$ $0.463$ $0.440$ $0.440$ $\boldsymbol{0.547}$* $0.378$ $0.503$ $0.428$ $0.412$ $0.430$ $\boldsymbol{0.492}$*
Answer in passage ($200$) $0.415$ $0.430$ $\boldsymbol{0.440}$ $0.410$ $0.400$ $\boldsymbol{0.440}$ $0.542$ $0.610$ $0.593$ $0.547$ $0.585$ $\boldsymbol{0.627}$*
Answer in list ($200$) $0.365$ $0.495$ $0.435$ $0.415$ $0.455$ $\boldsymbol{0.550}$* $0.355$ $0.507$ $0.410$ $0.453$ $0.438$ $\boldsymbol{0.613}$*
Answer in table ($200$) $0.540$ $0.555$ $0.540$ $0.520$ $0.495$ $\boldsymbol{0.580}$* $0.407$ $0.497$ $0.480$ $0.435$ $0.430$ $\boldsymbol{0.515}$*
English questions ($300$) $0.450$ $0.527$ $0.477$ $0.480$ $0.483$ $\boldsymbol{0.553}$* $0.423$ $0.525$ $0.475$ $0.465$ $0.475$ $\boldsymbol{0.573}$*
German questions ($300$) $0.430$ $0.460$ $0.467$ $0.417$ $0.417$ $\boldsymbol{0.493}$* $0.447$ $0.552$ $0.513$ $0.492$ $0.493$ $\boldsymbol{0.597}$*
Method. This simple idea is the core of our proposed Algorithm 1. The only other confounding factor is the presence of redundant evidences in the prompt: the retriever may sometimes fetch two (or more) semantically equivalent evidences if they both satisfy the user’s intent. In the context of counterfactuals, it may be that when one of these redundant evidences are removed, the answer hardly changes. This may lead to the wrong conclusion that the first evidence was unimportant: it just happens that the second evidence makes up for it content-wise. Hence, it is important that we first cluster evidences $E$ (Line $3$) by content before we apply an iterative removal of each evidence cluster $e^{cl}_{i}$ (Lines $4$ and $5$). Moreover, to adjust for non-determinism in LLM generations, we use a Monte Carlo method: in this case it simply means that we repeat the answer generation $m$ times (Lines $6-7$), compute similarities between the answer and the counterfactual answer in each iteration $a^{cf}_{i,j}$ (Line $8$), and average these $m$ values for each cluster (Line $10$). This average similarity is the contribution of each evidence cluster to the original answer. These raw answer contribution scores $s_{i}$ for each evidence cluster are then normalized via softmax to derive the final attribution distribution $\mathcal{A}$ (Fig. 4). We used the popular DBSCAN algorithm for clustering evidences (Line 3) as it requires only two tunable parameter ($\epsilon$ and $minPts$), and does not require the apriori specification of the number of clusters. Cosine similarity between the text embeddings of the answer and the counterfactual answer (in both cases the completed question from the benchmark is prepended for providing the context in which the similarity is sought) was used in Line 8, and using JinaAI embeddings [^4] https://huggingface.co/jinaai/jina-embeddings-v3 to encode evidences which was found to work better than other embedding models in this context. While our algorithm produces fairly accurate results (Table 3), output distributions (example in Fig. 4) may sometimes appear to be somewhat degenerate to end users. For example, say among six clusters, the top-evidence cluster is attributed $20\%$ and the remaining five groups are all assigned about $16\%$. We thus introduce a temperature parameter $t$ to skew the distribution more towards the top evidence cluster.
6. The ConfQuestions Benchmark
@@ 7. Results and Insights
Questions Attribution accuracy
All questions ($360$) $78.9\%$
Simple questions ($170$) $80.6\%$
Complex questions ($190$) $77.4\%$
Answer in passage ($94$) $77.7\%$
Answer in list ($121$) $82.6\%$
Answer in table ($145$) $76.6\%$
English questions ($190$) $77.9\%$
German questions ($170$) $80.0\%$
Table 3. Evaluation of explanation generation by RAGonite.
@@
Retr. P@1 $\boldsymbol{0.660}$ $0.550$ $0.470$ $0.420$ $0.460$ $0.580$
Ans. relevance $\boldsymbol{0.720}$ $0.515$ $0.560$ $0.495$ $0.585$ $0.635$
@@
Retr. P@1 $0.352$ $0.437$ $0.502$ $\boldsymbol{0.523}$
Ans. relevance $0.339$ $0.443$ $0.474$ $\boldsymbol{0.585}$
@@
Retr. P@1 $0.523$ $\boldsymbol{0.660}$
Ans. relevance $0.585$ $\boldsymbol{0.652}$
@@
Retr. P@1 $\boldsymbol{0.523}$ $0.502$ $0.497$ $0.302$ $0.515$
Ans. relevance $\boldsymbol{0.585}$ $0.561$ $0.565$ $0.408$ $0.573$
@@
Retr. P@1 $\boldsymbol{0.537}$ $0.505$ $0.523$
Ans. relevance $0.578$ $0.576$ $\boldsymbol{0.585}$
@@
Retr. P@1 $\boldsymbol{0.523}$ $0.485$
Ans. relevance $\boldsymbol{0.585}$ $0.517$
@@
Retr. P@1 $\boldsymbol{0.427}$ $0.403$
Ans. relevance $\boldsymbol{0.585}$ $0.561$
@@
Retr. P@1 $0.455$ $0.450$ $\boldsymbol{0.523}$
Ans. relevance $0.527$ $0.479$ $\boldsymbol{0.585}$
@@
Retr. P@1 $\boldsymbol{0.523}$ $0.427$
Ans. relevance $\boldsymbol{0.585}$ $\boldsymbol{0.585}$
@@ 7.1. Metrics
We use Precision@1 as our retrieval metric (since there is usually one gold URL), and Answer relevance as judged by gpt-4o (0 for non-relevant, 0.5 for partially relevant, 1 for relevant) against gold answers, as our metric for answer quality. Question-wise results are averaged over the full dataset. We use accuracy for evaluating attribution explanations as follows: we find the evidence that has the highest attribution score, and find its source URL. If this URL matches that of the gold answer, then accuracy is $1$, else 00. The average over all questions is then reported as a percentage.
@@ 7.2. Setup
We are now ready to run and evaluate RAGonite in various configurations over the 600 English and German questions in ConfQuestions. The proposed default configuration uses: (i) all available context; (ii) all evidences as the corpus; (iii) LLM-generated intent-explicit questions; (iv) verbalization as the linearization technique for table records; (v) both row- and table-level indexing for table elements; (vi) GPT-4o as our LLM of choice; (vii) BGE is as our text embedding model; (viii) hybrid search as our retrieval mode; and (ix) BGE reranking over the hybrid search results. Contextualization variants and attribution results are discussed in this section: all other ablation analysis is postponed to Sec. 7.4. For contextualization results, an asterisk (*) denotes statistically significant improvement of the proposed method over the strongest baseline of adding page titles for contextualization (second-best results in $5/8$ rows for Precision@1 and $/8$ 8 rows for Answer relevance in Table 2). We use the McNemar’s test for paired binomial data for Precision@1 since it is binary, and the 1-tailed paired $t$-test for Answer Relevance as it is non-binary. For a comparison to be marked as statistically significant, we use a $p$-value $<0.05$ for both cases.
@@ 7.3. Key findings
Evidence contextualization improves performance. Contextualization results are in Table 2. We make the following observations: (i) Adding all the suggested contexts (ALL) consistently led to significantly improved performances for retrieval and generation. (ii) Systematically, the page title (TTL) was found to be the single most helpful context to be added to the evidences, given the nature of our data here, with preceding headings (HDR) being the second best. Surrounding text (BEF and AFT) offered relatively lower gains in retrieval, but notably improved answer generation quality. (iii) These improvements are systematic w.r.t. question complexity (simple/complex), answer source (passage/list/table), and language (English/German), demonstrating RAGonite’s robust performance across diverse scenarios.
@@
Counterfactual attribution is effective. While counterfactual evidence has been proven to be a tool for discovering causal factors (Pearl and Mackenzie, 2020), our particular implementation still needed proof of efficacy. Results in Table 3 show that we consistently reach accuracies close to $80\%$, definitely an acceptable number in terms of user trust. The numbers in parentheses in Column 1 show over how many questions the respective averages were computed in Column 2. For questions where the gold document/evidence was not retrieved in the top-10, there is no chance that the attribution could spot the correct evidence. Hence these are removed from this evaluation and we do not see round numners like 600/200/100 as in Table 2. Looking at individual data slices over the rows, it is satisfying to observe that the high performance was not due to skewed success on easier slices, but is quite systematic over all criteria (varying only within a narrow span of $77-82\%$). Understandably, complex questions and cases where answers come from tables are slightly more difficult cases to solve. Performance for German was slightly better than that for English. Nevertheless, this current evaluation is rather coarse as we do not have gold annotations at evidence-level: a fine-grained study involving real users is left as future work.

### v2

@@ Abstract.
journalyear: 2025copyright: rightsretainedconference: Proceedings of the Eighteenth ACM International Conference on Web Search and Data Mining; March 10–14, 2025; Hannover, Germanybooktitle: Proceedings of the Eighteenth ACM International Conference on Web Search and Data Mining (WSDM ’25), March 10–14, 2025, Hannover, Germanydoi: 10.1145/3701551.3704126isbn: 979-8-4007-1329-3/25/03ccs: Information systems Question answering
@@ 1. Introduction
The benchmark and all other artifacts for this work are public at https://gitlab.cc-asp.fraunhofer.de/iis-nlp-public/ragonite-public and https://github.com/Fraunhofer-IIS/RAGonite.
@@ 2. System overview
Backend. An overview of the RAGonite pipeline is in Fig. 1. RAGonite’s backend is split into a functional core, which handles retrieval and answer generation, and a stateful layer that persists chats into a SQLite database and provides a REST API to the frontend using FastAPI. The dependencies in the functional core include the vector database (ChromaDB), and prompt template (Jinja) and LLM libraries (gpt-4o). We use ChromaDB as our vector database for storing the contextualized evidences, and also use ChromaDB’s in-built retrieval functions. The multilingual BGE embedding model [^1] https://huggingface.co/BAAI/bge-m3 was used to embed the documents, that was found to work slightly better than text-embedding-3-small [^2] https://openai.com/index/new-embedding-models-and-api-updates/ from OpenAI. While ChromaDB was used for its extensibility, we also explored variants like Weaviate and Milvus. We used top-k hybrid search with the BGE re-ranker [^3] https://huggingface.co/BAAI/bge-reranker-v2-m3 being used to merge $k$ dense retrieval and lexical retrieval results ($k=10$ in our experiments). We use gpt-4o as our LLM of choice inside the question completion and answer generation modules for efficiency and quality, but we also support other OpenAI, Mistral, and Llama models. All prompts can be seen inside the demo interface for transparency. A single GPU server (4x48GB NVIDIA Ada 6000 RTX, 512 GB RAM, 64 virtual cores) was used for all our experiments. All code is in Python.
@@ 2 Output: Distribution $\boldsymbol{\mathcal{A}}$ for attributing answer $a$ to each evidence $e$
Counterfactual Attribution (CFA). This simple idea is the core of our proposed Algorithm 1. The only other confounding factor is the presence of redundant evidences in the prompt: the retriever may sometimes fetch two (or more) semantically equivalent evidences if they both satisfy the user’s intent. In the context of counterfactuals, it may be that when one of these redundant evidences are removed, the answer hardly changes. This may lead to the wrong conclusion that the first evidence was unimportant: it just happens that the second evidence makes up for it content-wise. Hence, it is important that we first cluster evidences $E$ (Line $3$) by content before we apply an iterative removal of each evidence cluster $e^{cl}_{i}$ (Lines $4$ and $5$). Moreover, to adjust for non-determinism in LLM generations, we use a Monte Carlo method: in this case it simply means that we repeat the answer generation $m$ times (Lines $6-7$), compute similarities between the answer and the counterfactual answer in each iteration $a^{cf}_{i,j}$ (Line $8$), and average these $m$ values for each cluster (Line $10$). This average similarity is the contribution of each evidence cluster to the original answer. These raw answer contribution scores $s_{i}$ for each evidence cluster are then normalized via softmax to derive the final attribution distribution $\mathcal{A}$ (Fig. 4). We used the popular DBSCAN algorithm for clustering evidences (Line 3) as it requires only two tunable parameter ($\epsilon$ and $minPts$), and does not require the apriori specification of the number of clusters. Cosine similarity between the text embeddings of the answer and the counterfactual answer (in both cases the completed question from the benchmark is prepended for providing the context in which the similarity is sought) was used in Line 8, and using JinaAI embeddings [^4] https://huggingface.co/jinaai/jina-embeddings-v3 to encode evidences which was found to work better than other embedding models in this context. While our algorithm produces fairly accurate results (Table 3), output distributions (example in Fig. 4) may sometimes appear to be somewhat degenerate to end users. For example, say among six clusters, the top-evidence cluster is attributed $20\%$ and the remaining five groups are all assigned about $16\%$. We thus introduce a temperature parameter $t$ to skew the distribution more towards the top evidence cluster. Answers and counterfactual answers were generated using the cheaper gpt4o-mini model (instead of gpt4o) due to budget constraints: every question entails the generation of up to ten counterfactual answers, which entails a substantially higher number of output tokens than generating the intent-explicit questions or the original answers.
6. The ConfQuestions Benchmark
@@ 7. Results and Insights
Metric $\rightarrow$ Retrieval precision@1 Answer relevance
Data $\downarrow$ / Context $\rightarrow$ NONE +TTL +HDR +BEF +AFT +ALL NONE +TTL +HDR +BEF +AFT +ALL
All questions ($600$) $0.398$ $0.483$ $0.448$ $0.453$ $0.460$ $\boldsymbol{0.528}$ $0.388$ $0.477$ $0.435$ $0.445$ $0.404$ $\boldsymbol{0.529}$
Simple questions ($300$) $0.413$ $0.470$ $0.477$ $0.450$ $0.483$ $\boldsymbol{0.510}$ $0.423$ $0.537$ $0.482$ $0.502$ $0.458$ $\boldsymbol{0.593}$
Complex questions ($300$) $0.383$ $0.497$ $0.420$ $0.457$ $0.437$ $\boldsymbol{0.547}$ $0.352$ $0.417$ $0.388$ $0.388$ $0.350$ $\boldsymbol{0.465}$
Answer in passage ($200$) $0.365$ $0.410$ $0.390$ $0.415$ $0.420$ $\boldsymbol{0.445}$ $0.430$ $0.547$ $0.460$ $0.515$ $0.500$ $\boldsymbol{0.603}$
Answer in list ($200$) $0.340$ $0.475$ $0.440$ $0.435$ $0.490$ $\boldsymbol{0.560}$ $0.328$ $0.422$ $0.390$ $0.400$ $0.340$ $\boldsymbol{0.507}$
Answer in table ($200$) $0.490$ $0.565$ $0.515$ $0.510$ $0.470$ $\boldsymbol{0.580}$ $0.405$ $0.460$ $0.455$ $0.420$ $0.372$ $\boldsymbol{0.477}$
English questions ($300$) $0.413$ $0.500$ $0.467$ $0.480$ $0.483$ $\boldsymbol{0.563}$ $0.407$ $0.530$ $0.472$ $0.483$ $0.432$ $\boldsymbol{0.575}$
German questions ($300$) $0.383$ $0.467$ $0.430$ $0.427$ $0.437$ $\boldsymbol{0.493}$ $0.368$ $0.423$ $0.398$ $0.407$ $0.377$ $\boldsymbol{0.483}$
Data $\downarrow$ / Attribution $\rightarrow$ Naive CFA CFA w/ Clusters
All questions ($364$) $0.772$ $0.791$ $\boldsymbol{0.799}$
Simple questions ($175$) $0.771$ $0.811$ $\boldsymbol{0.817}$
Complex questions ($189$) $0.772$ $0.772$ $\boldsymbol{0.783}$
Answer in passage ($98$) $0.816$ $\boldsymbol{0.847}$ $0.806$
Answer in list ($123$) $0.756$ $0.780$ $\boldsymbol{0.821}$
Answer in table ($143$) $0.755$ $0.762$ $\boldsymbol{0.776}$
English questions ($192$) $\boldsymbol{0.807}$ $0.765$ $0.786$
German questions ($172$) $0.733$ $\boldsymbol{0.820}$ $0.814$
Table 3. Accuracies of explanation generation methods.
@@
Retr. P@1 $\boldsymbol{0.660}$ $0.530$ $0.500$ $0.430$ $0.490$ $0.560$
Ans. relevance $\boldsymbol{0.670}$ $0.495$ $0.465$ $0.440$ $0.555$ $0.550$
@@
Retr. P@1 $0.363$ $0.422$ $0.335$ $\boldsymbol{0.528}$
Ans. relevance $0.289$ $0.407$ $0.333$ $\boldsymbol{0.529}$
@@
Retr. P@1 $0.528$ $\boldsymbol{0.658}$
Ans. relevance $0.529$ $\boldsymbol{0.627}$
@@
Retr. P@1 $\boldsymbol{0.528}$ $0.392$ $0.382$ $0.368$ $0.372$
Ans. relevance $\boldsymbol{0.529}$ $0.364$ $0.363$ $0.366$ $0.361$
@@
Retr. P@1 $0.405$ $0.492$ $\boldsymbol{0.528}$
Ans. relevance $0.362$ $0.514$ $\boldsymbol{0.529}$
@@
Retr. P@1 $\boldsymbol{0.528}$ $0.480$
Ans. relevance $\boldsymbol{0.529}$ $0.449$
@@
Retr. P@1 $\boldsymbol{0.528}$ $0.525$
Ans. relevance $0.529$ $\boldsymbol{0.538}$
@@
Retr. P@1 $0.430$ $0.465$ $\boldsymbol{0.528}$
Ans. relevance $0.417$ $0.443$ $\boldsymbol{0.529}$
@@
Retr. P@1 $\boldsymbol{0.528}$ $0.417$
Ans. relevance $\boldsymbol{0.529}$ $0.507$
@@ 7.1. Metrics
We use Precision@1 as our retrieval metric (since there is usually one gold URL), and Answer relevance as judged by gpt-4o (00 for non-relevant, $0.5$ for partially relevant, $1$ for relevant) against gold answers, as our metric for answer quality. Question-wise results are averaged over the full dataset. We use accuracy for evaluating attribution explanations as follows: we find the evidence that has the highest attribution score, and find its source URL. If this URL matches that of the gold answer, then accuracy is $1$, else 00. The average accuracy over all questions is then reported, that is also interpretable as a percentage if multiplied by $100$.
@@ 7.2. Setup
We are now ready to run and evaluate RAGonite in various configurations over the $600$ English and German questions in ConfQuestions. The proposed default configuration uses: (i) all available context; (ii) all evidences as the corpus; (iii) LLM-generated intent-explicit questions; (iv) verbalization as the linearization technique for table records; (v) both row- and table-level indexing for table elements; (vi) GPT-4o as our LLM of choice; (vii) BGE is as our text embedding model; (viii) hybrid search as our retrieval mode; and (ix) BGE reranking over the hybrid search results. Contextualization variants and attribution results are discussed in this section: all other ablation analysis is postponed to Sec. 7.4.
Prompts. There are three prompts used in RAGonite: (i) for rephrasing a conversational question; (ii) for generating an answer from retrieved evidences; and (iii) for evaluating a generated answer. These are presented below in Figs. 67, and 8, respectively (used for both English and German questions).
Figure 6. LLM prompt for rephrasing a conversational question into an intent-explicit form.
Figure 7. LLM prompt for generating an answer from the retrieved pool of evidence.
Figure 8. LLM prompt for evaluating a generated answer according to the Answer Quality measure.
@@ 7.3. Key findings
Evidence contextualization improves performance. Contextualization results are in Table 2. We make the following observations: (i) Adding all the suggested contexts (ALL) consistently led to significantly improved performances for retrieval and generation. This means that enhancing evidences from documents with surrounding context is generally beneficial for helpful for RAG over collaborative wikis: content authors are often efficiency-oriented and do not repeat information inside passages, lists, or tables that are already mentioned in page titles or previous headings, for example. Similarly, adding preceding and succeeding texts like an overlapping chunking method, give coherence to the evidence and help in better retrieval (users often use localizing words in questions that may not be exactly inside the pertinent evidence but could be nearby in the document). Surrounding texts clearly also help the LLM generate more consistent and relevant responses; (ii) Systematically, the page title (TTL) was found to be the single most helpful context to be added to the evidences ($0.483$ P@1 for TTL vs. $0.528$ for ALL, and $0.477$ Answer Relevance for TTL vs. $0.529$ for ALL), with each of the other three individual add-ons brings in substantial improvements over NONE (ALL ¿ TTL ¿ AFT ¿ BEF ¿ HDR ¿ NONE for P@1 and ALL ¿ TTL ¿ BEF ¿ HDR ¿ AFT ¿ NONE for answer relevance). No feature hurts performance. It is notable that succeeding content also helps, as common coreferencing methods only contextualize using preceding text. (iii) These improvements are systematic with respect to question complexity (simple/complex), answer source (passage/list/table), and language (English/German), demonstrating why contextualization is a worthwhile operation across diverse question types. An interesting thing to note here that the biggest gains in retrieval ($0.220$ gain in P@1 of ALL over NONE) and answering ($0.179$ gain in answer relevance of ALL over NONE) performance are achieved for answer-in-list questions – a very common QA scenario for enterprise wikis where one asks for meeting notes, attendees and action items – all often stored as lists.
@@
Counterfactual attribution is effective. While counterfactual evidence has been proven to be a tool for discovering causal factors (Pearl and Mackenzie, 2020), our particular implementation still needed proof of efficacy. Results in Table 3 (Column $4$) show that we consistently reach accuracies close to $80\%$, definitely an acceptable number in terms of user trust. The numbers in parentheses in Column $1$ show over how many questions the respective averages were computed in Columns $2-4$. For questions where the gold document/evidence was not retrieved in the top-$10$, there is no chance that the attribution module could spot the correct evidence. Hence these are removed from this evaluation and we do not see round numbers like $600/200/100$ as in Table 2. Looking at individual data slices across rows, it is satisfying to observe that the high performance was not due to skewed success on easier slices: improvements over standard or naive attribution are quite systematic over almost all criteria (varying only within a narrow span of $77-82\%$). In the naive mode (also a baseline in (Cohen-Wang et al., 2024)), we compute cosine similarities between the answer and the evidences, and then use softmax scores over this distribution as the attribution scores for the corresponding evidences. The top scoring evidence is used to compute accuracy (Sec. 7.1). Within the counterfactual approaches, the one with clustering generally does better (with clusters best in $5/8$ cases, without clusters in $2/8$: marked in bold in Table 3). Understandably, complex questions and cases where answers come from tables are slightly more difficult cases to solve (relatively lower accuracies of $\simeq 77\%$). Performance for German was slightly better than that for English.